# Enhancing Edge Detection by Texture Handling Architecture and Noiseless Training Data

### v4

@@ 5.1.3 Preserving the Strongest Responses
Essentially, the distinction between edges and textures is based on response intensity to specific criteria, such as gradients. However, textural regions can also produce high responses, leading to false positives. SDPED addresses this issue through the use of CSDB, which can be viewed as an averaging mechanism across multiple SDB responses. This design reinforces pixels consistently identified as edges across multiple layers while suppressing less reliable responses. If a pixel exhibits a high response in only a single SDB layer but low responses in others, its final response is attenuated due to the skipping connections. As a result, SDPED effectively preserves only the most consistent edge responses while minimizing the influence of textural noise.
5.1.4 Strengthening Feature Fusion
ED architectures can be broadly categorized into feature extraction and feature fusion stages. While previous models allocate tens of millions of parameters to multi-scale feature extraction, they often rely on a single $1\times 1$ convolution layer for feature fusion. This stark imbalance between extraction and fusion limits the full utilization of extracted features. To address this limitation, SDPED incorporates an extended feature fusion block with additional layers, enhancing the integration of information and improving ED performance.
5.2 Model Size of SDPED
As shown in Table 4, SDPED has the fewest parameters among all compared models. Its model size is less than half that of previous architectures, which is particularly noteworthy given that it achieves comparable or superior performance. This result suggests that the SDPED architecture is more parameter-efficient, utilizing model weights more effectively than prior models. In contrast, the significantly larger number of parameters in previous models may indicate redundancy or inefficiencies in their design.
@@ 5.2 Model Size of SDPED
Size (kB) Num. of Coe. (×$\times 10^{6}$)
HED (2015) 57499 14.716171
RCF (2017) 57850 14.803781
RIND (2021) 232593 59.388526
BDCN (2022) 63749 16.302120
Dexi (2023) 137775 35.215245
SDPED_$\_7$(Ours) 22470 5.728638
5.3 Effectiveness Across Different ED Datasets
One of the key strengths of SDPED is its robustness across multiple ED datasets. As highlighted in [39], previous models often demonstrate dataset-specific performance, excelling on certain datasets while underperforming on others. For instance, HED and RCF achieve the highest accuracy on BRIND but are outperformed on UDED and BIPED2. Similarly, BDCN performs best on UDED but lags on other datasets, while Dexi and RIND yield superior results on MDBD and BIPED2 but fail to maintain consistency across different benchmarks.
In contrast, SDPED consistently achieves SOTA performance across various datasets. It yields the best results in all benchmarks on BRIND, UDED, and MDBD and significantly outperforms previous models in AP on BIPED2 while maintaining competitive performance on other benchmarks. This demonstrates its adaptability across different data distributions. These results indicate that SDPED is not only highly precise but also versatile, making it a robust solution for ED tasks across diverse datasets.
@@ 5.4 Limitations and Future Work
Despite the effectiveness of our methods, there remain areas for further improvement.
5.4.1 Inference Speed and Memory Consumption
Since SDPED processes full-resolution images, it requires more GPU memory and may have slower inference speeds compared to previous models. In this study, our primary focus was to enhance ED performance rather than optimize computational efficiency. However, in practical applications, a trade-off between accuracy and computational cost is often necessary. Future research could explore acceleration strategies and memory-efficient implementations to enhance real-time applicability.
5.4.2 Feature Fusion Schemes
Although ablation studies have shown the effectiveness of the extended feature fusion block, merely adding extra convolution layers may not be the optimal strategy for further improving edge detection performance. Since feature fusion has been largely overlooked in ED tasks, exploring more advanced fusion schemes could yield additional performance gains.
5.4.3 Noiseless ED Data
A fundamental challenge in ED tasks is the lack of a formal definition for edges. If edges had a strict mathematical definition, traditional algorithms would have already solved ED problems. As a result, ED datasets heavily rely on human-labeled annotations, which inherently introduce noise.
As demonstrated in Section IV, model performance is often correlated with dataset quality. For instance, all models perform best on UDED, despite it having the fewest training samples, likely due to its superior annotation quality. This suggests that data quality may be a more significant bottleneck than model architecture in ED tasks.
While Section III introduces an initial method for obtaining noiseless data, further research is needed to develop larger, high-quality datasets with minimal noise.
@@ 6 Conclusion
In this paper, we proposed SDPED, a novel class of precise ED models, and introduced a data augmentation strategy to improve model performance on noiseless data. Our model achieves SOTA performance across multiple datasets and partitions, particularly demonstrating significant improvements in AP. The results confirm that SDPED is both more precise and more adaptable than previous models. Furthermore, SDPED achieves superior performance with fewer parameters, utilizing less than half the coefficients of previous SOTA models. This highlights the parameter efficiency and architectural effectiveness of SDPED. Finally, our data augmentation strategy provides the first approach for leveraging noiseless data in training ED models, opening new possibilities for future research in obtaining high-quality ED data.

### v5

@@ 5.1.3 Consistency-based Response Filtering
Essentially, the distinction between edges and textures is based on response intensity to specific criteria, such as gradients, but textural regions can also produce high responses, leading to false positives. Here, the CSDB structure acts as an implicit consistency-filtering mechanism. Because it cascades multiple SDB blocks with skipping connections, the final prediction effectively reinforces pixels that are consistently identified as edges across multiple processing stages. If a textural pixel produces a high gradient response in only a single layer but lacks consistency in subsequent blocks, its final response is attenuated through the integration mechanism. This design encourages the preservation of more reliable edge responses, effectively filtering out sporadic textural noise that often plagues shallower or non-cascaded models.
5.1.4 Restoring Structural Balance in Feature Fusion
ED architectures can be broadly categorized into feature extraction and feature fusion stages. However, there is a significant structural imbalance in previous ED models: millions of parameters are often dedicated to feature extraction, yet the critical fusion stage is reduced to a single $1\times 1$ convolution layer. This bottleneck may lead to the underutilization of rich extracted features. By incorporating a more robust fusing block with three convolutional layers (including $3\times 3$), SDPED restores the balance between extraction and fusion, allowing for a more sophisticated synthesis of edge features and resulting in sharper, more accurate final edge maps.
5.2 Parameter Efficiency and Model Size
As summarized in Table 7, SDPED exhibits strong parameter efficiency. It achieves SOTA performance with the fewest parameters among all compared models—less than 40% of the second-smallest model (HED) and only 1.0% of the largest (EdgeNat). This suggests that SDPED utilizes its parameters more efficiently than previous backbones, namely, with less redundancy to achieve high-precision edge localization.
@@ 5.2 Parameter Efficiency and Model Size
Size (kB) Num. of Coe. ($\times 10^{6}$)
HED 57499 14.716171
RCF 57850 14.803781
RIND 232593 59.388526
BDCN 63749 16.302120
Dexi 137775 35.215245
EdgeNat 2349905 587.34319
SDPED_7\_7 (Ours) 22470 5.728638
5.3 Robustness and Cross-Dataset Adaptability
One of the key strengths of SDPED is its robustness across multiple ED datasets. As highlighted in [42], previous models often exhibit dataset-dependent performance, with varying behavior across benchmarks. For instance, HED and RCF achieve the highest accuracy on BRIND but are outperformed on UDED and BIPED2. Similarly, BDCN performs best on UDED but lags on other datasets, while Dexi and RIND yield superior results on MDBD and BIPED2 but fail to maintain consistency across different benchmarks.
In contrast, SDPED demonstrates consistently competitive, and often state-of-the-art, performance across various datasets. It yields the best results in all benchmarks on BRIND, UDED, and MDBD and significantly outperforms previous models in AP on BIPED2 while maintaining competitive performance on other benchmarks. This demonstrates its adaptability across different data distributions. These results indicate that SDPED is not only highly precise but also versatile, making it a robust solution for ED tasks across diverse datasets.
@@ 5.4 Limitations and Future Work
While our framework sets a new benchmark for high-precision ED, several avenues for future work remain:
5.4.1 Inference Efficiency and Memory Footprint
The primary trade-offs for SDPED’s high precision are its memory consumption and inference speed. Since the model operates on full-resolution images without down-sampling, it requires more GPU memory and inference time compared to models that utilize pooling. While this study prioritized precision over memory and speed, future work should explore memory-efficient implementations to adapt SDPED for applications on resource-constrained devices.
5.4.2 Advanced Fusion Mechanisms
Although our multi-layer fusing block improves performance, as confirmed by ablation studies, it remains a relatively straightforward convolutional integration. Given that feature fusion could be the bottleneck in high-precision tasks, exploring more diverse and advanced schemes—such as attention-based fusion or dynamic feature selection—could further refine the ability of leveraging extracted rich edge information.
5.4.3 The Data Quality Bottleneck
A fundamental challenge in ED tasks is the lack of a formal definition for edges. If edges had a strict mathematical definition, traditional algorithms would have already solved ED problems. As a result, ED datasets heavily rely on human-labeled annotations, which inherently introduce noise. As demonstrated in Section 4, model performance is often correlated with dataset quality. For instance, all models perform best on UDED, despite it having the fewest training samples, likely due to its superior annotation quality. This suggests that data quality may represent a bottleneck comparable to model architecture in ED tasks. While our ideal prior method in Section 3.2 provides a way to leverage noiseless data, the development of meticulously curated datasets and data-refinement approaches remains essential for further performance breakthroughs in ED.
@@ 6 Conclusion
In this paper, we have presented a comprehensive framework for high-precision edge detection (ED). By performing a task-adaptive re-engineering of texture-handling structures originally developed for image super-resolution (SR), the proposed SDPED model, built upon the Cascaded Skipping Density Block (CSDB), is able to suppress textural responses while preserving high spatial fidelity. Extensive experimental results across multiple benchmarks demonstrate that SDPED achieves strong performance, particularly in Average Precision (AP), while maintaining high parameter efficiency. These results indicate that a precisely tailored and compact architecture can be more effective than substantially larger models, and that certain architectural backbones exhibit meaningful task transferability across low-level vision problems. Beyond architectural advancements, we introduce an ideal prior guidance strategy—a data-centric augmentation paradigm that incorporates noiseless ground-truth information into the training process. By treating labels as noise-free input samples, this approach provides a practical mechanism to bypass the human-annotation ceiling, which inevitably introduces subjectivity and noise into supervision. More broadly, this perspective suggests that ground-truth information can be exploited as an ideal physical prior to guide learning beyond the limitations imposed by noisy supervision, providing a potential solution for a wide range of learning tasks, particularly for those where human annotation is unavoidable. Finally, we address inconsistencies in ED model assessment by analyzing and adopting a 1-pixel evaluation criterion. By shifting evaluation from resolution-dependent proportional ratios to a fixed-pixel standard, this criterion offers a more rigorous and resolution-independent benchmark for high-precision ED. Taken together, this work provides a coherent study of high-precision edge detection from architectural, data-centric, and evaluation perspectives. We hope these insights contribute to the development of more reliable and precise edge detectors and inspire further research on precision-oriented modeling in low-level and soft computing–based vision tasks.